OpenStudy (anonymous):

multiply (3√5-1)(6√5-4)

OpenStudy (ghazi):

\[(3 \sqrt5 -1)(6 \sqrt5 -4)=18 \sqrt5 * \sqrt5 -12 \sqrt 5-6 \sqrt5 +4 =18*5 -18 \sqrt 5 +4 \]

OpenStudy (ghazi):

94- 18 sqrt5

OpenStudy (anonymous):

thank you so much!

OpenStudy (ghazi):

:)